Staff:

LBTO: Guido, Brandon,Xianju, Juan Carlos

Summary.

Run the z0 and housekeeper patch for a goal of 10000 iterations at frequency 1500Hz.

The adsec is setup for th test version. The close loop at 1500 showed to be unstable and the same for the 1350Hz. To proceed with the acquisition of the write z0 (ncpa) statistics a 1000 iterations could be completed.

The time allowed to work until 12:00 sharp.

Details.

8:00 am power on.

Argos cal unit reported not possible deploy fom OSS.

State: retracted.

Status: legal.

swim arm

park position yes

in working no

locked no

hardware status: simotion Error

absolutte encoder 56.78

Status

enable codition no

index switch#1 no

switch#2 activated no

telescope yes, enable bymcs yes,enabled by argos yes, PFC in park positions yes

RESET DRIVER.

ENABLE CONDITIONS FULLFILLED: YES.

power on
  • ocam not live.
  • power cycle bcu2k
  • restart configuraiton.

test plan:

https://docs.google.com/spreadsheets/d/1val68NAFIEzbmn0hsACgCK3RrXHdlt84hT-k5HPQbWI/edit#gid=0

  • Adsec switch to test conf.
  • AOS connected
  • wind speed not available.
  • restart MOXA from internet service.
  • set flat.
  • 8:53 modes #641

change AO table for operations:

use table: table_LUCIFER_ACE-AO.txt.20221012_testz0

test close loop 1500

  • bad control. modes not good saved TN: 20221012_164954
  • many actuators in warning for rms.
  • Xianyu checking the adsec telemetry.
  • The oscillaion of the peak is observed in the modes and residual modes TT and higher modes.
modalplot_20221012_164954.png

plotjitter_20221012_164954.png mode0_psd_20221012-164954.png

mode1_psd_20221012_164954.png mode2_psd_20221012_164954.png

mode3_psd_20221012_164954.png mode4_psd_20221012_164954.png

residualmodes0_psd_20221012_164954.png residualmodes3_psd_20221012_164954.png

9:55 am Adsec rest/flat.

change ao table to setup freq at 13500

FREQ 1350

  • Loop unstable shows to be similar to the previous one.
  • _171426
modalplot_20221012_171426.png plotjitter_20221012_171426.png

switch to 990

20221012_173322

9.5 4.20E4 1 1500 500 KL_v20 20190930_161000 iir1350HzV001 20181103_001410 3 4 2 0 1 4 4.3 20 800 200 100 30 10e-9 80.0

10.5 1.67E4 1 1500 600 KL_v20 20190930_161000 iir1350HzV001 20181103_001410 3 4 1 0 1 4 4.3 20 800 200 300 30 10e-9 80.0

11.4 7.30E3 1 1500 600 KL_v20 20190930_161000 iir1350HzV001 20181103_001410 3 4 1 0 1 16 12.6 20 800 200 600 30 10e-9 80.0

11.4 7.30E3 1 1500 600 KL_v20 20190930_161000 iir1350HzV001 20181103_001410 3 4
  • launch script.
  • python run-ncpa.py --count 10000 --delay 3 --bank 1
  • Brandon fixing bugs in the script.
  • TN:20221012_184545

11:55 am stop

  • sate safe.
  • rest shell
  • change to use_soul stable.

-- JuanCarlosGuerra - 12 Oct 2022
Topic revision: r4 - 13 Oct 2022, JuanCarlosGuerra
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback